It is known that the mass of the earth is 81 times the mass of the moon.Show that the point of weightless between the earth and the moon for a spacecraft occur at 9/10 of the distance to the moon Hint: Me/b²=Mm/(a-b)²?


1
Expert's answer
2022-03-08T08:55:29-0500

At the equilibrium point


"G\frac{M_em}{b^2}=G\frac{M_mm}{(a-b)^2}"

Hence


"\frac{a}{b}-1=\sqrt{\frac{M_m}{M_e}}=9""\frac{a}{b}=10""b=1\/10a"

Here "a" is the distance from Earth to the Moon.

Finally, the distance from earth to the equilibrium point is


"a-b=a-1\/10a=9\/10a"